Good day

The Ksp is 5.61 x10^-11
From the Ksp,I calculated:
The molar solubility of Mg(OH)2 in pure water is 2.41 x 10 ^ -4 M
The molar solubility of Mg(OH)2 In 0.100 M NaOH is 5.61 x 10^-9
Please calculate the ratio of solubility of Mg(OH)2 dissolved in pure H2O to Mg(OH)2 dissolved in a 0.100 M NaOH solution.
Express answer to 3 significant figures please

Thank you so much.
1
Expert's answer
2020-08-05T05:06:36-0400

Solution.

"С_1(Mg(OH)_2)=2.41\sdot10^{-4}M;"

"C_2(Mg(OH)_2)=5.61\sdot10^{-9}M;"

"\dfrac{C_1}{C_2}=\dfrac{2.41\sdot10^{-4}M}{5.61\sdot10^{-9}M}=4.30\sdot10^{4};"


Answer: "\dfrac{C_1}{C_2}=4.30\sdot10^{4}."